# Large-charge Rényi entropy

### v1

@@ I Introduction
In this paper, we compute information theoretic quantities in strongly-coupled CFTs using the large-charge expansion. This will result in one of the first concrete, holography-free computations of quantum-information-theoretic quantities in strongly-coupled CFTs. The quantity of our interest is the charged vacuum Rényi entanglement entropy on a disk, denoted as $S_{n}(Q)$. It is defined for theories with a global symmetry (we will hereafter focus on $U(1)$) as the Rényi entropy of the density matrix $\tilde{\rho}_{A,Q}\equiv\rho_{A}\mathcal{P}_{Q}/\mathop{\rm Tr}\nolimits\left[\rho_{A}\mathcal{P}_{Q}\right]$, where $\mathcal{P}_{Q}$ is a projection operator onto the charge-$Q$ subspace and $\rho_{A}$ is the reduced density matrix on region $A$ [5, 6]. The quantity has interesting applications such as quantifying the speed of symmetry restoration after a symmetry-breaking quantum quench [7, 8, 9, 10, 11], where our result is potentially useful.
@@
Aside from being the first such computations, we will find several interesting and universal results. For one thing, we find that the probability of realising charge-$Q$ vacuum fluctuations inside a disk scales as $\exp\left(-Q^{3/2}\right)$ at large-$Q$, for any three-dimensional CFTs with a proper thermodynamic limit. We also find that $S_{n}(Q)$ converges to a universal $O(1)$ number at large-$Q$, for any CFTs described by the same large-charge EFT, called the large-charge universality class hereafter [2]. As a concrete example, such a number is computed for the large-charge universality class including the $O(2)$ Wilson-Fisher CFT.
@@
The rest of the paper is organised as follows. We start in Section II with a short account of the charged Rényi entropy, in particular its relation to the charged thermal partition function on a hyperbolic spatial slice. We will also discuss at length the regularisation we use in this paper. We will then give a brief review of the large-charge expansion in Section III, using which we compute the charged Rényi entropy for a broad class of CFTs in Section IV. We conclude in Section V with a summary of results and the outlook.
II Charged Rényi entropy
II.1 Charge-resolved density matrix
@@
The charged Rényi entanglement entropy is a variant of the entanglement entropy for systems with a global symmetry, taken to be $U(1)$ hereafter for simplicity [5, 6]. Let the reduced density matrix of a (uncharged) state restricted to a region $A$ be $\rho_{A}$. In the presence of a symmetry, it can be decomposed into a direct sum of smaller matrices in each charged sector,
$\rho_{A}=\bigoplus_{Q}\rho_{A,Q}.$ (1)
@@
Note that each $\rho_{A,Q}$ is not a density matrix per se, because $p(Q)\equiv\mathop{\rm Tr}\nolimits\left[\rho_{A,Q}\right]\neq 1$. Rather, $p(Q)$ gives an overall probability of realising charge-$Q$ fluctuations in region $A$. On the other hand, the charged $n$-th Rényi entanglement entropy $S_{n}(Q)$ is related to the normalised reduced density matrix proportional to $\rho_{A,Q}$, defined as
$S_{n}(Q)\equiv\frac{1}{1-n}\log\left(\frac{\mathop{\rm Tr}\nolimits\left[\rho_{A,Q}^{n}\right]}{\mathop{\rm Tr}\nolimits\left[\rho_{A,Q}\right]^{n}}\right).$ (2)
From now on, we will implicitly take the region $A$ as a disk of radius $R$ in $\mathbb{R}^{D-1}$, and the state in the total Hilbert space as vacuum, |$\ket{0}$. For later convenience we will take the radial coordinates $\mathbb{R}^{D-1}$ whose origin is at the centre of the disk, with the radial direction denoted as $r$.
II.2 A conformal map to the hyperbolic disk
@@
As shown in [12], the modular Hamiltonian on a disk of a CFT is given by the time-evolution operator of the same CFT on $H^{D-1}$. In other words, we have $\rho_{A}={e^{-2\pi RH_{\mathrm{m}}}}/{\mathop{\rm Tr}\nolimits\left[e^{-2\pi RH_{\mathrm{m}}}\right]}$, where $H_{\mathrm{m}}$ is the Hamiltonian of the CFT defined on $H^{D-1}$. Likewise, $\rho_{A,Q}$ can be written as
$\rho_{A,Q}=\frac{e^{-2\pi RH_{\mathrm{m}}}}{\mathop{\rm Tr}\nolimits\left[e^{-2\pi RH_{\mathrm{m}}}\right]}\mathcal{P}_{Q}.$ (3)
@@
using the projection operator on the charge-$Q$ subspace, $\mathcal{P}_{Q}$. By plugging this into (2), we see that the charged Rényi entropy is given by
$S_{n}(Q)\equiv\frac{\log Z_{n}(Q)-n\log Z_{1}(Q)}{1-n},$ (4)
@@
where $Z_{n}(Q)\equiv\mathop{\rm Tr}\nolimits\left[e^{-2\pi RnH_{\mathrm{m}}}\mathcal{P}_{Q}\right]$. In other words, $S_{n}(Q)$ can be obtained from the charge-$Q$ partition function on a thermal geometry $S^{1}\times H^{D-1}$ with metric
$ds^{2}$ $=d\tau^{2}+R^{2}\left(du^{2}+\sinh^{2}u\,d\Omega_{D-2}^{2}\right),$ (5)
@@
II.3 UV regularisation and the boundary condition
@@
The correct regularisation is as follows: modify the original state with a factorisation map that separates the subregion and its compliment physically with boundary conditions, $\mathcal{M}_{\delta}:\,H\to H_{A}\otimes H_{B}$. [^1] For more about the definition of tensor factorisation of the CFT Hilbert space used to define the entanglement entropy and the importance of boundary conditions there, see [13, 14, 15]. Let us, to simplify the argument, take $H_{A}$ and $H_{B}$ to be the CFT Hilbert spaces with conformal boundary conditions imposed on the surfaces defined by $r=R-\delta$ and $r=R\left(1-\delta/R\right)^{-1}$, respectively. [^2] As we will see later, this corresponds to placing the boundary condition on a constant $u$ slice in the hyperbolic frame. We will defer specifying the boundary condition until after conformal transforming the geometry to $S^{1}\times H^{D-1}$. By using such a map, the regularised reduced density matrix is defined as $\rho_{A,\delta}\equiv\mathop{\rm Tr}\nolimits_{B}\left[M_{\delta}\rho M^{\dagger}_{\delta}\right]$, and this has discrete spectrum as promised for compact CFTs.
@@
$u=\log\left(\frac{2R}{\delta}+1\right)\equiv u_{0}.$ (6)
@@
$V_{2}=2\pi\left(\frac{e^{u_{0}}}{2}-1+\frac{e^{-u_{0}}}{2}\right)$ (7)
$Z(\beta,Q)\equiv\mathop{\rm Tr}\nolimits\left[e^{-\beta H_{\mathrm{m}}^{\mathrm{reg}}}\mathcal{P}_{Q}\right].$ (8)
III Large-charge expansion
A crucial observation in the large-charge expansion is that physical quantities involving a large-charge ground state |$\ket{Q}$ can be computed by using a saddle-point configuration which is helical in time, order-by-order in the $1/Q$-expansion. Such a VEV will look like $\phi=Ae^{i\mu t}$ ($t$ is the Lorentzian time in (5), $t=-i\tau$) in the Wilson-Fisher case, and so the $O(2)$ symmetry (as well as the conformal symmetry) is spontaneously broken. Note that this is consistent with the Neumann boundary condition we have imposed. [^3] See [16] for more about the effect of boundary conditions at large global charge. We are therefore entitled to write down an EFT using the Nambu-Goldstone (NG) boson $\chi$, where we have decomposed the field as $\phi=ae^{i\chi}$.
@@ III Large-charge expansion
$L=\tilde{c}_{3/2}\left|\partial\chi\right|^{3}+\tilde{c}_{1/2}{\tt Ric}_{3}\left|\partial\chi\right|+O(\rho^{-1/2}),$ (9)
[…]
IV Large-charge Rényi entropy
IV.1 Large-charge spectrum
Let us start by deriving the dispersion relation on top of our large-charge vacuum, $\chi=\mu t$, which will serve as a starting point for deriving the spectrum at large-$Q$ and then computing the thermal partition function. By separating $\chi$ into VEV and fluctuation, $\chi=\mu t+\hat{\chi}$, the fluctuation Lagrangian at leading order, which is $O(1)$, becomes
$L\ni\hat{\chi}\left(-\partial_{t}^{2}+\frac{1}{2R^{2}}\triangle_{H^{2}_{\rm reg}}\right)\hat{\chi},$ (10)
@@ IV.1 Large-charge spectrum
$\omega_{m,k}^{2}R^{2}=\frac{1}{2}\left(\frac{1}{4}+s^{2}_{m,k}\right)$ (11)
where $\frac{1}{4}+s^{2}_{m,k}$ denotes the $k$-th eigenvalue of $\triangle_{H^{2}_{\rm reg}}$ in the sector of angular momentum $m$, and the corresponding eigenfunction is $\chi_{m,k}(u,\theta)\equiv P^{m}_{-1/2+is_{m,k}}(\cosh u)e^{im(\theta+i\pi)}\in\mathbb{R}$. As already implicit, this one-particle spectrum is discrete instead of continuous because of the Neumann boundary condition.
Because the discrete spectrum is hard to obtain analytically, we have computed $\omega_{m,k}$ numerically for various $e^{u_{0}}\gg 1$. This is in anticipation that the final result will be organised in terms of $e^{u_{0}}$, which will be justified a posteriori when we compute the Rényi entropy. We also truncated the computation at $s_{m,k}=3$ because each mode labelled by $s_{m,k}\gtrsim 3$ will contribute to the Rényi entropy at order $O(e^{-2\pi\omega_{m,k}})=O(10^{-6})$. We show an example of the spectrum for $u_{0}=3$ in Table 1.
Table 1: List of $s_{m,k}\leq 3$ subject to the Neumann boundary condition at $u_{0}=3$.
@@
${{E_{\rm g.s.}(Q)}R}={V_{2}}\left(c_{3/2}\,q^{\frac{3}{2}}+c_{1/2}\,q^{\frac{1}{2}}+\alpha+O(q^{-\frac{1}{2}})\right),$ (12)
where $q\equiv Q/V_{2}$. Note that we are taking $Q\to\infty$ limit first, while taking $V_{2}$ to be large but fixed in terms of $Q$ (see Section II.3). While $c_{*}$ are the Wilsonian coefficients and cannot be determined from the EFT analysis alone, the one-loop contribution $\alpha$ is a number that is in principle calculable. While it is universal among theories in the same large-charge universality class and hence is in general important [2, 19], its precise value is of no use to us later on – we leave its computation for future work.
@@
IV.2 Large-charge fluctuation probability
Let us compute $p(Q)\equiv\mathop{\rm Tr}\nolimits\left[\rho_{A,Q}\right]$ which gives the total probability of measuring charge-$Q$ fluctuations on a disk of radius $R$ in vacuum. We can use (3) to rewrite $p(Q)$ as
$p(Q)=\frac{\mathop{\rm Tr}\nolimits\left[e^{-2\pi RH_{\mathrm{m}}}\mathcal{P}_{Q}\right]}{\mathop{\rm Tr}\nolimits\left[e^{-2\pi RH_{\mathrm{m}}}\right]}=\frac{Z(2\pi R,Q)}{Z(2\pi R)},$ (13)

### v2

@@ 1 Introduction
In this paper, we compute information theoretic quantities in strongly-coupled CFTs using the large-charge expansion. This will result in one of the first concrete, holography-free computations of quantum-information-theoretic quantities in strongly-coupled CFTs. The quantity of our interest is the charged vacuum Rényi entanglement entropy on a disk, denoted as $S_{n}(Q)$. It is defined for theories with a global symmetry (we will hereafter focus on $U(1)$) as the Rényi entropy of the density matrix $\tilde{\rho}_{A,Q}\equiv\rho_{A}\mathcal{P}_{Q}/{\mathrm{Tr}}\left[\rho_{A}\mathcal{P}_{Q}\right]$, where $\mathcal{P}_{Q}$ is a projection operator onto the charge-$Q$ subspace and $\rho_{A}$ is the reduced density matrix on region $A$ [5, 6]. The quantity has interesting applications such as quantifying the speed of symmetry restoration after a symmetry-breaking quantum quench [7, 8, 9, 10, 11], where our result is potentially useful.
@@
Aside from being the first such computations, we will find several interesting and universal results. For one thing, we find that the probability of realising charge-$Q$ vacuum fluctuations inside a disk scales as $e^{-Q^{3/2}}$ at large-$Q$, for any three-dimensional CFTs with a proper thermodynamic limit. We also find that $S_{n}(Q)$ converges to a universal $O(1)$ number at large-$Q$, for any CFTs described by the same large-charge EFT, called the large-charge universality class hereafter [2]. As a concrete example, such a number is computed for the large-charge universality class including the $O(2)$ Wilson-Fisher CFT.
@@
The rest of the paper is organised as follows. We start in Section 2 with a short account of the charged Rényi entropy, in particular its relation to the charged thermal partition function on a hyperbolic spatial slice. We will also discuss at length the regularisation we use in this paper. We will then give a brief review of the large-charge expansion in Section 3, using which we compute the charged Rényi entropy for a broad class of CFTs in Section 4. We conclude in Section 5 with a summary of results and the outlook.
2 Charged Rényi entropy
2.1 Charge-resolved density matrix
@@
The charged Rényi entanglement entropy is a variant of the entanglement entropy for systems with a global symmetry, taken to be $U(1)$ hereafter for simplicity [5, 6, 13]. Let the reduced density matrix of a (uncharged) state restricted to a region $A$ be $\rho_{A}$. In the presence of a symmetry, it can be decomposed into a direct sum of smaller matrices in each charged sector,
$\rho_{A}=\bigoplus_{Q}\rho_{A,Q}.$ (2.1)
@@
Note that each $\rho_{A,Q}$ is not a density matrix per se, because $p(Q)\equiv{\mathrm{Tr}}\left[\rho_{A,Q}\right]\neq 1$. Rather, $p(Q)$ gives an overall probability of realising charge-$Q$ fluctuations in region $A$. On the other hand, the charged $n$-th Rényi entanglement entropy $S_{n}(Q)$ is related to the normalised reduced density matrix proportional to $\rho_{A,Q}$, defined as
$S_{n}(Q)\equiv\frac{1}{1-n}\log\left(\frac{{\mathrm{Tr}}\left[\rho_{A,Q}^{n}\right]}{{\mathrm{Tr}}\left[\rho_{A,Q}\right]^{n}}\right).$ (2.2)
[…]
2.2 A conformal map to the hyperbolic disk
@@
As shown in [12], the modular Hamiltonian on a disk of a CFT is given by the time-evolution operator of the same CFT on $H^{D-1}$. In other words, we have $\rho_{A}={e^{-2\pi RH_{\mathrm{m}}}}/{{\mathrm{Tr}}\left[e^{-2\pi RH_{\mathrm{m}}}\right]}$, where $H_{\mathrm{m}}$ is the Hamiltonian of the CFT defined on $H^{D-1}$. Likewise, $\rho_{A,Q}$ can be written as
$\rho_{A,Q}=\frac{e^{-2\pi RH_{\mathrm{m}}}}{{\mathrm{Tr}}\left[e^{-2\pi RH_{\mathrm{m}}}\right]}\mathcal{P}_{Q}.$ (2.3)
@@
using the projection operator on the charge-$Q$ subspace, $\mathcal{P}_{Q}$. By plugging this into (2.2), we see that the charged Rényi entropy is given by
$S_{n}(Q)\equiv\frac{\log Z_{n}(Q)-n\log Z_{1}(Q)}{1-n},$ (2.4)
@@
where $Z_{n}(Q)\equiv{\mathrm{Tr}}\left[e^{-2\pi RnH_{\mathrm{m}}}\mathcal{P}_{Q}\right]$. In other words, $S_{n}(Q)$ can be obtained from the charge-$Q$ partition function on a thermal geometry $S^{1}\times H^{D-1}$ with metric
$ds^{2}$ $=d\tau^{2}+R^{2}\left(du^{2}+\sinh^{2}u\,d\Omega_{D-2}^{2}\right),$ (2.5)
@@
2.3 UV regularisation and the boundary condition
@@
The correct regularisation is as follows: modify the original state with a factorisation map that separates the subregion and its compliment physically with boundary conditions, $\mathcal{M}_{\delta}:\,H\to H_{A}\otimes H_{B}$. [^2] The factorisation scale $\delta$ is a physical scale which separates the subregion from its compliment, and can be thought of as different from the UV cutoff below which the theory is conformal. In particular, it makes complete sense to have modular energy as large as one wants, after fixing $\delta$. For more about the definition of tensor factorisation of the CFT Hilbert space used to define the entanglement entropy and the importance of boundary conditions there, see [14, 15, 16]. Let us, to simplify the argument, take $H_{A}$ and $H_{B}$ to be the CFT Hilbert spaces with conformal boundary conditions imposed on the surfaces defined by $r=R-\delta$ and $r=R\left(1-\delta/R\right)^{-1}$, respectively. [^3] As we will see later, this corresponds to placing the boundary condition on a constant $u$ slice in the hyperbolic frame. We will defer specifying the boundary condition until after conformal transforming the geometry to $S^{1}\times H^{D-1}$. By using such a map, the regularised reduced density matrix is defined as $\rho_{A,\delta}\equiv{\mathrm{Tr}}_{B}\left[M_{\delta}\rho M^{\dagger}_{\delta}\right]$, and this has discrete spectrum as promised for compact CFTs.
@@
$u=\log\left(\frac{2R}{\delta}+1\right)\equiv u_{0}.$ (2.6)
@@
The regularisation fixes the volume of the regularised hyperbolic space to be $\mathop{{\tt Vol}}(H^{D-1}_{\rm reg})=\mathop{{\tt Vol}}(S^{D-2}_{\rm reg})\times\int_{0}^{u_{0}}du\,\sinh^{D-2}(u)$, where $\mathop{{\tt Vol}}(S^{D-2}_{\rm reg})$ is the area of the (D2)(D-2)-sphere of radius $R$. For later convenience we define a rescaled volume, ${V}_{D-1}\equiv\mathop{{\tt Vol}}(H^{D-1}_{\rm reg})/R^{D-1}$; We will focus on $D=3$ hereafter, in which case we have
$V_{2}=2\pi\left(\frac{e^{u_{0}}}{2}-1+\frac{e^{-u_{0}}}{2}\right)$ (2.7)
$Z(\beta,Q)\equiv{\mathrm{Tr}}\left[e^{-\beta H_{\mathrm{m}}^{\mathrm{reg}}}\mathcal{P}_{Q}\right].$ (2.8)
3 Large-charge expansion
@@
$L=\tilde{c}_{3/2}\left|\partial\chi\right|^{3}+\tilde{c}_{1/2}{\tt Ric}_{3}\left|\partial\chi\right|+O(\rho^{-1/2}),$ (3.1)
where $\left|\partial\chi\right|\equiv\sqrt{-\partial_{\mu}\chi\partial^{\mu}\chi}$, $\tt Ric$} is the Ricci scalar of the underlying geometry (which is a constant $(D-1)(D-2)/R^{2}$ for $\mathbb{R}\times H^{D-1}$), and $\tilde{c}_{3/2,\,1/2}$ are some Wilsonian coefficients. Note that this action is to be evaluated around a linear background, $\chi=\mu t$. The charge density, given as $\rho\equiv{\frac{Q}{\mathop{{\tt Vol}}(H^{2}_{\rm reg})}}$, is related to $\mu$ via the Noether’s theorem so that $\mu=O\left(\sqrt{\rho}\right)$. The subscripts of $c_{*}$ therefore denote the scaling at which the terms contribute, e.g., $\tilde{c}_{3/2}$ is the Wilsonian coefficient at $O(\rho^{3/2})$.
Leaving coefficients such as $\tilde{c}_{3/2,\,1/2}$ undetermined is an unfortunate feature of the effective action, but it has its own benefits. Take any theory with the $O(2)$ symmetry; if we know that the only light mode around the large-charge saddle-point is the NG boson, it is then immediate, because of the underlying conformal and $O(2)$ symmetries, that the same effective action (with different Wilsonian coefficients) describes the large-charge sector of that theory. We can express this by saying that theories with the same large-charge EFT define a large-charge universality class. For example, (3.1) describes not only the $O(2)$ Wilson-Fisher CFT at large charge, but the $O(N)$ Wilson-Fisher CFT at large-rank symmetric traceless representations [18] and $\mathbb{C}P^{N-1}$ model at large monopole number [19].
4 Large-charge Rényi entropy
4.1 Dispersion relation
Let us start by deriving the dispersion relation on top of our large-charge vacuum, $\chi=\mu t$. This will serve as a starting point for computing the thermal partition function $Z(\beta,Q)$, as the resulting one-particle spectrum governs the leading contributions to the energy spectrum in the large-$Q$ limit (see (4.5), for example.) By separating $\chi$ into VEV and fluctuation, $\chi=\mu t+\hat{\chi}$, the fluctuation Lagrangian at leading order, which is $O(1)$, becomes
$L\ni\hat{\chi}\left(-\partial_{t}^{2}+\frac{1}{2R^{2}}\triangle_{H^{2}_{\rm reg}}\right)\hat{\chi}$, (4.1)
@@ 4.1 Dispersion relation
$\omega_{m,k}^{2}R^{2}=\frac{1}{2}\left(\frac{1}{4}+s^{2}_{m,k}\right)$ (4.2)
where $\frac{1}{4}+s^{2}_{m,k}$ denotes the $k$-th eigenvalue of $\triangle_{H^{2}_{\rm reg}}$ in the sector of angular momentum $m$, and the corresponding eigenfunction is $\chi_{m,k}(u,\theta)\equiv P^{m}_{-1/2+is_{m,k}}(\cosh u)e^{im\theta+i\pi}\in\mathbb{R}$. As already implicit, this one-particle spectrum is discrete instead of continuous because of the Neumann boundary condition. [^5] One might think that the continuous spectrum suitably approximates the discrete spectrum, but that would be too quick. Because the IR cutoff is always kept finite, at large enough angular momentum $m$, the continuous description is never correct. In fact we would get a wrong coefficients for $A_{n}$ had we used the continuous approximation. It is extremely important that the spectrum, at leading order in large-$Q$, does not depend on the Wilsonian coefficients, $c_{*}$, at all. This fact will translate to the universality of the charged Rènyi entropy at large-$Q$, as shown in (4.8).
Because the discrete spectrum is hard to obtain analytically, we have computed $\omega_{m,k}$ numerically for various $e^{u_{0}}\gg 1$. [^6] They were obtained by locating the zeros of $\partial_{u}\chi_{m,k}(u,\theta)$ by using a bisection method. See https://github.com/machiparu/lcre for the Python program. This is in anticipation that the final result will be organised in terms of $e^{u_{0}}$, which will be justified a posteriori when we compute the Rényi entropy. We also truncated the computation at $s_{m,k}=3$ because each mode labelled by $s_{m,k}\gtrsim 3$ will contribute to the Rényi entropy at order $O(e^{-2\pi\omega_{m,k}})=O(10^{-6})$. We show an example of the spectrum for $u_{0}=3$ in Table 1.
Table 1: List of $s_{m,k}\leq 3$, which gives the one-particle spectrum subject to the Neumann boundary condition via (4.2), at $u_{0}=3$.
@@ 4.2 Spectra at large-charge
${{E_{\rm g.s.}(Q)}R}={V_{2}}\left(c_{3/2}\,q^{\frac{3}{2}}+c_{1/2}\,q^{\frac{1}{2}}+\alpha+O(q^{-\frac{1}{2}})\right),$ (4.3)
where $q\equiv Q/V_{2}$. Note that we are taking $Q\to\infty$ limit first, while taking $V_{2}$ to be large but fixed in terms of $Q$ (see Section 2.3). While $c_{*}$ are the Wilsonian coefficients and cannot be determined from the EFT analysis alone, the one-loop contribution $\alpha$ is a number that is in principle calculable. [^7] For the $O(2)$ Wilson-Fisher CFT, one can use the operator dimensions of $\phi^{Q}$ obtained via the Monte-Carlo simulation [20, 21, 22] to read out $(\tilde{c}_{3/2},\tilde{c}_{1/2})=(0.103,-0.020)$. This translates to $({c}_{3/2},{c}_{1/2})=(1.20,-0.370)$. While it is universal among theories in the same large-charge universality class and hence is in general important [2, 23], its precise value is of no use to us later on – we leave its computation for future work.
@@
4.3 Large-charge fluctuation probability
Let us compute $p(Q)\equiv{\mathrm{Tr}}\left[\rho_{A,Q}\right]$ which gives the total probability of measuring charge-$Q$ fluctuations on a disk of radius $R$ in vacuum. We can use (2.3) to rewrite $p(Q)$ as
$p(Q)=\frac{{\mathrm{Tr}}\left[e^{-2\pi RH_{\mathrm{m}}}\mathcal{P}_{Q}\right]}{{\mathrm{Tr}}\left[e^{-2\pi RH_{\mathrm{m}}}\right]}=\frac{Z(2\pi R,Q)}{Z(2\pi R)},$ (4.4)